2. Suppose country B is a small export country (price taker) in the world market of lamp trade. Its domestic demand on and supply for lamp are: D=18-P, S=-12+2P. The world market price of lamp is 14.

(1)  Calculate the equilibrium price and quantity of lamp in country B before trade.

(2)  Calculate B’s quantity of exported lamp under free trade.

Draw graph and calculate the net welfare effect of free trade on B’s consumers, producers and the whole country.


1
Expert's answer
2021-09-12T19:31:28-0400

a)

"QP=18-P\\QS=-12+2P\\Qd=Qs\\18-P=-12+2P\\18+12=3P\\P=10\\Q=18-10\\Q=8units"


b)

World price 14

Required domestic

"Q's=-12+2(14)=16units\\Qd=18-14=4units\\Qs>Qd, surplus"

There will be export of "16-4=12units"


c)

"Consumer\space surplus\\0=18-P\\P=18\\\frac{1}{2}×(18-14)×4\\=\$8"

"Producer\space surplus\\0=-12+2P\\2P=12\\P=6\\\frac{1}{2}×(14-6)×16\\=\$64"



.